Fermi National Laboratory
 

Operation Summary for July 25, 2011 - July 27, 2011

 

Monday July 25

The day shift began with NuMI off due to correction magnet (V108) cabling problems, with Operations monitoring the Tevatron (TEV) store 8927 (152.5E30) and the Recycler stash (158.7E10), with the Antiproton Source (Pbar) stacking (11.2E10), with the Main Injector (MI), Linac, and Booster operating normally, with the MiniBooNE and Meson FTBFT-1008 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, a MiniBooNE person asked to have their beam turned off at 8:47 AM for a MI-12 service building access to look at the air conditioning (AC) unit. Operators reported that technicians were investigating problems with the Recycler stochastic cooling system. The Cryo Coordinator reported that the Pbar dry engine had a slipping belt. Cryo system personnel began repairs at 10:20 AM and finished them about 40 minutes later. Operators reported at 2:15 PM that they saw a problem with the TEV proton and antiproton tunes and made adjustments; Unfortunately, their adjustments went a little too far and caused a small loss of luminosity. Operators corrected their mistake. At 2:29 PM, store 8927 quenched due to a CDF beam abort. At 2:31 PM, the Pbar antiproton abort kicker (C:PAK1) suffered a prefire. Kicker Personnel began repairs. MiniBooNE completed their work and Operations resumed sending them beam at 2:56 PM. Operators prepared D0 for an access at 3:30 PM.

During the evening shift, Operations reported at 4:58 PM that the Recycler momentum stochastic cooling system was off and should remain off. D0 came out of access at 5:26 PM. MTA - MuCool personnel reported at 5:29 PM that they had completed an hour of studies. At 7:49 PM, operators reported that the TEV quenched during ramping. A TEV person came in and reported at 9:25 PM that the quench was due to a D0 quadrupole magnet (D0Q3). Operators reported that the TEV sector A0 pond water pump was off. FESS personnel investigated and found the pump motor had grounded. With the help of security and the Fire Department, they ran a hose from the fire hydrant, across the street to the compressors as a temporary supply for their cooling water.

Tuesday July 26

The midnight shift began with the TEV off for sector A0 pond water work, with Operations monitoring the stash (397.4E10), with Pbar stacking (119.3E10), and with MiniBooNE taking beam.

During the midnight shift, Operations disabled beam to T-1008 at 12:16 AM. The A0 hose installation was completed by 2:51 AM. Operators put the TEV through a dry squeeze at 3:11 AM. Operators put the TEV through a wet squeeze at 4:05 AM. Operations put the TEV into shot setup at 4:28 AM. Operators turned off the Booster beam at 5:15 AM in preparation for an access on the next shift; operators put Pbar into standby and turned off beam to MI. Operations established store 8929 at 5:27 AM with an average initial luminosity of 390.8E30. Water Works personnel began cleaning the heat exchanger at F1 at 6:51 AM. Linac technicians reported at 7:46 AM that they were beginning their work on a RF station (LRF2); the Linac was off for this work.

During the day shift, operators prepared MI for access at 8:04 AM. Operators prepared MI-8 and MI-12A & B for access at 8:18 AM. Operators prepared MI-65 for access at 8:28 AM. EE Support and contract electrical personnel reported at 9:44 AM that the cables running down from the service building were intact. The shorted cables were in the short penetration between the MI side and the NuMI stub in the tunnel. Linac personnel reported that they would be turning off all their Klystrons to switch from LCW water skid #2 to #1 to repair a leak. Booster work ended and Booster resumed sending beam at 12:14 PM. Technicians repaired a Booster RF station (BRF13) LCW leak with epoxy; the station will remain off until Wednesday to allow the epoxy to cure. At 3:16 PM, Linac personnel reported that LRF3 needed its power amplifier (PA) replaced.

During the evening shift, operators reported at 10:01 PM that they had searched and secured MI and had MI ramping by 11:07 PM. Operations established beam to the Linac dump, to MiniBooNE, and beam to the MI dump by 11:33 PM. Linac personnel reported that they had changed the PA for LRF3 twice today: the first PA showed a anode to grid short after they installed it.

Wednesday July 27

The midnight shift began with Operations monitoring store 8929 (28.51E30) and the stash (157.3E10), with Pbar stacking (3.3E10), and with MiniBooNE taking beam.

During the midnight shift, Operations established beam to MiniBooNE at 12:06 AM. A Pbar person began work on a Pbar Debuncher RF station (DRF2) at 12:20 AM; this will lower the stacking rate. Operators prepared Pbar for access at 12:52 AM. The Pbar person reported at 2:29 AM that the DRF2 cavity fan exhaust vent was heavily clogged, which was causing fan problems. Pbar resumed stacking at 2:37 AM. Operations terminated store 8929 at 3:51 AM, put the TEV into shot setup, and established store 8930 at 5:25 AM with an AIL of 164.1E30. Operators prepared Booster and MI for access at 6:09 AM.

The Plans for Wednesday

The plans for today are to monitor store 8930 while the NuMI V108 cabling work continues. The Run Coordinator will allow MTA to take beam and for TEV personnel to conduct a crystal collimator end-of-store study.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/27/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory